OpenStudy (anonymous):

For exponential functions of the form f(x) = a^x which of the following expressions is equal to f(1) x f (2) ? answer choices: A: f(1^2) B: f(1 x 2) C: f(1 +2)

OpenStudy (anonymous):

Note that \[ f(1) \cdot f(2) = a^1 \cdot a^2 = a^3 = a^{1+2} = f(1+2), \]so the answer is C.